# jina-clip-v2: Multilingual Multimodal Embeddings for Text and Images

### v1

@@ 4.1 Cross-modal Evaluation
We evaluate jina-clip-v2 on a set of English and multilingual cross-modal tasks, conducting a comparative analysis against both its predecessor jina-clip-v1 and the NLLB-CLIP-SigLIP (Visheratin, 2023b) large and base variants, a top-performing multilingual CLIP model. Additionally, we track the model’s performance at different training stages to evaluate the effectiveness of different parts of the training protocol. Measuring 865M parameters, our model sits between the two NLLB-CLIP-SigLIP versions in terms of size: nllb-siglip-base (507M parameters, 41% smaller than ours) and nllb-siglip-large (1.2B parameters, 39% larger than ours).
For English zero-shot image-text and text-image retrieval, we conduct evaluations on Flickr30K (Young et al., 2014) and MS COCO Captions (Chen et al., 2015), both of which are included in the CLIP Benchmark. For multilingual cross-modal tasks, we assess performance on the Crossmodal-3600 (Thapliyal et al., 2022) and multilingual MS COCO (Chen et al., 2015) datasets.
As shown in Table 2, jina-clip-v2 demonstrates strong performance across both English and multilingual benchmarks while showing clear progression through its training stages. On the English CLIP Benchmark, the final model achieves 79.09% and 89.73% on text-to-image and image-to-text retrieval respectively, showing significant improvement from stage 1’s 73.87% and 86.61%. While nllb-siglip-large leads in text-to-image retrieval at 81.54%, jina-clip-v2 outperforms its predecessor jina-clip-v1 and surpasses both NLLB-CLIP-SigLIP variants in image-to-text retrieval by 2.1% and 1.6% respectively. On the multilingual front, we obtain competitive results on Crossmodal-3600, achieving 81.43% and 83.23% on text-to-image and image-to-text retrieval respectively, approaching the performance of nllb-siglip-large. On multilingual MS COCO (Chen et al., 2015), jina-clip-v2 performs comparably to NLLB-CLIP-SigLIP models, reaching 84.87% and 86.03% on text-to-image and image-to-text retrieval. The improvement from Stage 1 to the final model is notable across all benchmarks, with gains of up to 10.6 percentage points on Crossmodal-3600’s text-to-image retrieval, demonstrating the effectiveness of the training strategy.
@@ 4.2 Text Retrieval and STS Evaluation
| Benchmark | Retrieval | STS | ||
| Metric | nDCG@10 [%] | Spearman Correlation [%] | ||
| Model - Language | EN | Multilingual | EN | Multilingual |
| nllb-siglip-large | 24.91 | - | 74.89 | - |
| jina-embeddings-v3 | 53.87 | 72.58 | 85.80 | 69.81 |
| jina-clip-v1 | 48.33 | - | 80.92 | - |
| jina-clip-v2stage 1 | 40.91 | - | 79.67 | - |
| jina-clip-v2stage 2 | 43.17 | - | 80.33 | - |
| jina-clip-v2 | 49.32 | 69.85 | 81.29 | 67.77 |
Table 3 presents the evaluation on MTEB tasks, demonstrating the effectiveness of jina-clip-v2’s text encoder for both retrieval and semantic textual similarity (STS) tasks. On English retrieval tasks (nDCG@10), jina-clip-v2 achieves an average score of 49.32, a significant improvement over nllb-siglip-large’s 24.91 and a slight improvement over jina-clip-v1’s 48.33, though not reaching jina-embeddings-v3’s score of 53.87. For multilingual retrieval, the model achieves 69.85, approaching jina-embeddings-v3’s performance of 72.58. In STS tasks measured by Spearman Correlation, jina-clip-v2 demonstrates competitive performance with 81.29 for English tasks, surpassing nllb-siglip-large’s 74.89 and comparable to jina-clip-v1’s 80.92, though below jina-embeddings-v3’s 85.8. On multilingual STS tasks, the model achieves 67.77, slightly below jina-embeddings-v3’s 69.8.
@@ 4.3 Visual Document Retrieval
| Model | Average (nDCG@5) [%] |
| jina-clip-v1 | 17.72 |
| nllb-siglip-large | 46.55 |
| jina-clip-v2 | 52.65 |
| jina-clip-v2stage 1 | 37.37 |
| jina-clip-v2stage 2 | 47.76 |
Visual document retrieval requires vision models to capture fine-grained information from images to accurately match documents with queries. This task is particularly challenging as it deals with high-resolution document images, which differ significantly from typical image-caption datasets. As shown in table 4, jina-clip-v2 achieves state-of-the-art performance on the ViDoRe Benchmark (Faysse et al., 2024) for visual document understanding, with an average nDCG@5 score of 52.65%, significantly outperforming both its predecessor jina-clip-v1 (17.72%) and nllb-siglip-large (46.55%). The progression through training stages demonstrates substantial improvements, advancing from 37.37% in stage 1 to 47.76% in stage 2, before reaching peak performance in the final model. These results highlight the effectiveness of our training strategy and the impact of the new datasets introduced for document retrieval tasks.
@@ 4.4 Matryoshka Representation Learning
| Avg CLIP Benchmark | 79.10 | 79.12 | 78.93 | 78.32 | 75.90 | 70.51 | ||||||
| Avg Crossmodal-3600 | 81.43 | 82.35 | 82.31 | 81.75 | 78.17 | 72.52 | ||||||
| Avg Multilingual MS COCO | 84.87 | 84.85 | 84.60 | 84.32 | 81.80 | 77.85 | ||||||
@@
| Avg CLIP Benchmark | 89.73 | 89.60 | 89.55 | 89.35 | 87.48 | 83.20 | ||||||
| Avg Crossmodal-3600 | 83.23 | 83.26 | 83.21 | 82.81 | 80.54 | 75.37 | ||||||
| Avg Multilingual MS COCO | 86.03 | 86.02 | 86.02 | 85.84 | 84.37 | 81.02 | ||||||
@@
| Avg STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
Table 5 presents the impact of embedding truncation on the performance of our model. The table is divided into 4 task categories: text-to-image retrieval, image-to-text retrieval, text-to-text retrieval and semantic textual similarity (STS).
The results show that performance remains highly stable when reducing dimensions from 1024 to 256, with minimal degradation (typically less than 1%) across all evaluation tasks. At 256 dimensions, representing a 75% reduction in embedding size, the model effectively preserves essential semantic information. Substantial performance degradation is only observed at very low dimensions (128 and 64), indicating the model’s very robust acquisition of image and text features.
@@ 5 Analysis
5.1 The Role of Image Resolution in Complex Document Retrieval
@@
Our empirical study demonstrates that image resolution plays a critical role in retrieving visually rich documents. To better understand the relationship between image resolution and model performance on such documents, we have conducted a targeted experiment.
@@
We checkpointed jina-clip-v2 at the end of Stage 1, when it had only seen images with a resolution of (224, 224). This model checkpoint was trained on 1.45 billion image-caption pairs and 1.45 billion text-text pairs. We conducted four runs, keeping the image resolution to for the first run and increasing to , , and for three additional runs. Each run is trained for 3500 steps using the same visually rich training set and hyperparameters. The models are then evaluated on the ViDoRe benchmark (Faysse et al., 2024), which comprises 10 datasets designed for retrieving visually rich documents, based on text queries.
@@
We have plotted the results in Figure 2. Unsurprisingly, increasing image resolution has a positive impact on linking captions to visually rich documents. The most significant improvement occurs when resolution increases from to , with the average nDCG@5 score across 10 benchmarks rising from to . A further increase in resolution to also brings a noticeable gain in performance.
@@
Considering that the model is trained with a patch size of 14, increasing the resolution beyond significantly raises the number of patches, a increase when moving from to . This leads to substantial additional computational overhead while yielding only marginal improvement, with nDCG@5 increasing by just . Therefore, we identify as an optimal image resolution, offering a good balance between performance and computational efficiency.
5.2 Unified Batch vs Multi-Task Learning
@@
Our training approach, as outlined in Section 3, involves two tasks at each stage, both optimized in a multi-task learning framework: a multi-modal alignment task for images and text, and second text-only task. Both use the InfoNCE (Van den Oord et al., 2018) objective. Regardless of whether the loss is computed on image-text or text-text pairs, the same loss function is applied to pairs of vectors. This raises the question: Would unifying the two batches of pairs into a single batch, irrespective of modality, and optimizing a single contrastive objective, offer any benefits?
@@
We refer to this approach as the unified batch technique, in contrast to the multi-task learning paradigm. We present the training paradigm in Figure 3. The motivation is two-fold: first, simplicity of computation and secondly, modality gap mitigation (Liang et al., 2022) by using in-batch negatives across modalities. Compared to multi-task training, the unified batch approach utilizes a single InfoNCE loss with a shared temperature value, attempting to force all modalities to align within the same embedding space. Additionally, we integrate a SimCLR-like self-supervised learning method (Chen et al., 2020) into the contrastive training, aiming to provide more robust image representation.
@@
However, our Stage 1 experiments show that this technique does bring some early improvements, but fails beyond a certain point, at least with regard to cross-modal tasks. This was reflected in the slow decay of the loss temperature (our temperature is a trainable parameter in this case) and its eventual stabilization around . This is in contrast to the multi-task learning approach, where the temperature value decreases more rapidly and keeps decreasing throughout training, leading to better performance on cross-modal tasks.
@@
We hypothesize that this limitation arises from the fundamental information asymmetry between the visual and textual modalities (Schrodi et al., 2024). Images include fine-grained details of information whereas the corresponding textual descriptions include more coarse-grained level of information (Schrodi et al., 2024; Chen et al., 2023b). In temperature-scaled contrastive learning, the temperature parameter controls the hardness of the loss function with respect to the negative samples, i.e., the level of penalties on the hard negative samples (Wang & Liu, 2021). From an information-theory perspective, the level of penalty on the hard negative text-text samples theoretically differs from that of image-text samples due to the information gap problem. Consequently, as supported by our experiments, enforcing a unified temperature parameter across both modalities is potentially theoretically suboptimal.
@@ 6 Conclusion
In this work we present an enhanced strategy for training dual-encoder vision-language embedding models for cross-modal and text retrieval tasks, using contrastive learning. We introduce improvements to the pipeline, namely support for multiple languages, Matryoshka Representation Learning, and finer visual perception. We have trained and released a new model, jina-clip-v2, which has strong cross-modal and text retrieval performance on standard benchmarks. Finally, we have conducted analyses of three important considerations for CLIP models going forward: the ability to understand complex visual inputs performance how the image resolution affects it, and the effects of efforts to mitigate the modality gap.
@@ Appendix A Appendix
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||
| average | 81.43 | 73.51 | 84.13 | 82.07 | 79.29 | |||||||||
| ar | 73.56 | 66.22 | 76.89 | 78.92 | 76.94 | |||||||||
| bn | 63.78 | 50.19 | 68.11 | 75.19 | 74.19 | |||||||||
| da | 85.39 | 76.53 | 87.97 | 87.14 | 86.64 | |||||||||
| de | 91.25 | 84.64 | 92.42 | 89.56 | 87.25 | |||||||||
| el | 75.03 | 68.53 | 77.92 | 77.83 | 71.97 | |||||||||
| en | 75.83 | 69.78 | 77.78 | 73.11 | 72.22 | |||||||||
| es | 83.64 | 78.28 | 86.28 | 82.64 | 79.97 | |||||||||
| fi | 82.83 | 75.28 | 85.89 | 86.42 | 81.44 | |||||||||
| fr | 88.78 | 83.50 | 90.67 | 87.86 | 85.58 | |||||||||
| hi | 55.25 | 42.03 | 59.47 | 60.31 | 58.08 | |||||||||
| id | 84.22 | 73.69 | 86.94 | 86.31 | 84.17 | |||||||||
| it | 88.33 | 81.19 | 89.64 | 85.94 | 82.67 | |||||||||
| ja | 87.03 | 77.28 | 90.03 | 86.06 | 83.14 | |||||||||
| ko | 78.81 | 71.81 | 83.22 | 78.75 | 75.47 | |||||||||
| nl | 82.56 | 75.72 | 84.47 | 81.69 | 78.86 | |||||||||
| no | 81.08 | 71.97 | 83.08 | 82.69 | 79.97 | |||||||||
| pl | 84.00 | 79.33 | 86.50 | 82.72 | 78.61 | |||||||||
| pt | 82.42 | 76.17 | 85.19 | 82.69 | 79.44 | |||||||||
| ro | 89.36 | 82.92 | 92.22 | 90.03 | 86.17 | |||||||||
| ru | 88.97 | 82.97 | 91.11 | 86.44 | 83.78 | |||||||||
| sv | 78.06 | 71.33 | 80.56 | 79.33 | 76.17 | |||||||||
| th | 81.61 | 70.92 | 85.08 | 81.14 | 78.83 | |||||||||
| tr | 81.31 | 75.31 | 84.53 | 83.47 | 81.00 | |||||||||
| uk | 88.56 | 82.28 | 89.89 | 85.44 | 81.89 | |||||||||
| vi | 86.64 | 76.31 | 89.06 | 85.56 | 82.56 | |||||||||
| zh | 78.97 | 66.97 | 82.50 | 76.56 | 74.64 | |||||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||
| average | 83.23 | 79.64 | 84.12 | 80.16 | 76.56 | |||||||||
| ar | 76.25 | 72.14 | 76.67 | 75.86 | 73.69 | |||||||||
| bn | 69.00 | 63.78 | 70.22 | 75.58 | 73.61 | |||||||||
| da | 88.53 | 84.42 | 88.86 | 86.53 | 84.69 | |||||||||
| de | 92.47 | 88.42 | 92.47 | 87.50 | 84.44 | |||||||||
| el | 73.33 | 73.22 | 75.61 | 74.81 | 68.89 | |||||||||
| en | 78.58 | 74.33 | 79.61 | 70.81 | 69.08 | |||||||||
| es | 86.28 | 81.19 | 87.36 | 81.19 | 77.00 | |||||||||
| fi | 84.19 | 81.17 | 85.64 | 84.25 | 78.75 | |||||||||
| fr | 90.89 | 86.56 | 91.00 | 86.64 | 83.64 | |||||||||
| hi | 61.64 | 54.39 | 61.56 | 61.89 | 59.83 | |||||||||
| id | 86.31 | 83.64 | 87.64 | 84.33 | 81.89 | |||||||||
| it | 90.17 | 84.31 | 90.53 | 83.50 | 78.64 | |||||||||
| ja | 88.50 | 85.53 | 89.44 | 84.03 | 81.14 | |||||||||
| ko | 81.42 | 79.22 | 83.00 | 76.75 | 74.17 | |||||||||
| nl | 82.47 | 77.94 | 83.33 | 79.28 | 74.33 | |||||||||
| no | 83.75 | 77.22 | 84.42 | 82.08 | 77.03 | |||||||||
| pl | 84.61 | 83.72 | 85.69 | 79.58 | 75.39 | |||||||||
| pt | 83.94 | 79.31 | 84.06 | 79.39 | 74.89 | |||||||||
| ro | 91.31 | 87.39 | 92.08 | 88.83 | 83.58 | |||||||||
| ru | 90.64 | 87.75 | 90.89 | 84.64 | 80.61 | |||||||||
| sv | 78.28 | 77.53 | 80.11 | 76.58 | 72.31 | |||||||||
| th | 81.94 | 80.22 | 84.11 | 79.56 | 76.28 | |||||||||
| tr | 82.67 | 80.36 | 83.81 | 80.36 | 77.69 | |||||||||
| uk | 89.53 | 86.42 | 89.97 | 83.22 | 78.67 | |||||||||
| vi | 88.06 | 85.64 | 88.58 | 83.14 | 80.03 | |||||||||
| zh | 79.22 | 74.94 | 80.42 | 73.83 | 70.22 | |||||||||

### v2

@@ 4.1 Crossmodal Retrieval Evaluation
We evaluate jina-clip-v2 on a set of English and multilingual crossmodal tasks, conducting a comparative analysis with both jina-clip-v1 and the state-of-the-art NLLB-CLIP (Visheratin, 2023b) large and base SigLIP variants. Additionally, we track the model’s performance at all training stages to evaluate the effectiveness of the training protocol. For English zero-shot image-text and text-image retrieval, we conduct evaluations on the CLIP Benchmark, which includes Flickr30K (Young et al., 2014) and COCO Captions (Chen et al., 2015). For multilingual crossmodal retrieval tasks, we assess performance on the Crossmodal-3600 (Thapliyal et al., 2022) and XTD10 (Aggarwal & Kale, 2020; Rajendran et al., 2016) datasets.
Table 2 demonstrates the strong performance of jina-clip-v2 across both English and multilingual benchmarks. On the English CLIP Benchmark, our model outperforms jina-clip-v1 and while nllb-siglip-large leads in text-to-image retrieval, jina-clip-v2 outperforms both NLLB-CLIP variants in image-to-text retrieval. On multilingual multimodal retrieval, we obtain competitive results, approaching the performance of nllb-siglip-large in text-to-image retrieval and surpassing it in image-to-text retrieval. Regarding stage progression, performance improves considerably between Stage 1 and Stage 2 on all benchmarks, but drops slightly from Stage 2 to Stage 3. We hypothesize that there is a trade-off between crossmodal and text-only retrieval performance, and during stage 3 training, where focus shifts towards text embedding training with hard negatives, the model falls slightly behind on crossmodal retrieval.
@@ 4.2 Text Retrieval and STS Evaluation
| Retrieval nDCG@10 [%] | |||||||||
| Model | #Parameters∗ | en | zh | hi | de | fr | es | jp | ru |
| jina-clip-v1 | 137M | 47.04 | 8.09 | 8.73 | 33.77 | 39.99 | 39.77 | 10.77 | 4.78 |
| nllb-siglip-base | 414M | 24.69 | 27.04 | 39.18 | 22.67 | 28.30 | 30.06 | 30.73 | 32.68 |
| nllb-siglip-large | 767M | 27.83 | 32.96 | 43.22 | 32.29 | 38.74 | 39.92 | 35.53 | 38.87 |
| jina-embeddings-v3 | 572M | 48.34 | 64.27 | 65.18 | 63.59 | 64.09 | 62.97 | 68.47 | 66.46 |
| jina-clip-v2stage 1 | 561M | 42.02 | 55.11 | 62.03 | 57.56 | 58.56 | 59.86 | 63.34 | 60.74 |
| jina-clip-v2stage 2 | 561M | 43.47 | 57.66 | 62.19 | 59.16 | 61.05 | 61.51 | 65.13 | 63.66 |
| jina-clip-v2 | 561M | 46.46 | 60.47 | 63.08 | 56.66 | 61.16 | 61.48 | 65.48 | 64.73 |
| STS Spearman correlation based on cosine similarity | |||||||||
| Model | #Parameters∗ | en | zh | hi | de | fr | es | jp | ru |
| jina-clip-v1 | 137M | 81.35 | 22.97 | 36.09 | 50.04 | 67.97 | 70.05 | 55.83 | 39.77 |
| nllb-siglip-base | 414M | 72.57 | 33.36 | 68.11 | 39.58 | 63.94 | 65.28 | 77.39 | 57.55 |
| nllb-siglip-large | 767M | 74.73 | 39.15 | 73.78 | 53.45 | 72.38 | 69.68 | 79.25 | 62.84 |
| jina-embeddings-v3 | 572M | 81.68 | 54.60 | 83.62 | 75.32 | 81.11 | 80.78 | 81.03 | 77.16 |
| jina-clip-v2stage 1 | 561M | 80.01 | 52.74 | 83.35 | 74.84 | 79.06 | 80.42 | 81.48 | 75.62 |
| jina-clip-v2stage 2 | 561M | 80.76 | 53.37 | 84.30 | 75.64 | 80.19 | 80.51 | 81.46 | 76.18 |
| jina-clip-v2 | 561M | 81.58 | 55.08 | 79.55 | 76.27 | 81.70 | 81.89 | 82.01 | 77.79 |
∗ Refers to the parameters of the text tower
Table 3 presents the results of evaluations on the retrieval and semantic textual similarity (STS) tasks from MTEB (Muennighoff et al., 2023), both in English-only and multilingual contexts. jina-clip-v2 significantly improves on the performance of nllb-siglip-large in both English and multilingual text tasks, demonstrating the value of explicit text retrieval training. Compared to jina-clip-v1, performance is significantly better on multilingual tasks, and comparable on English-only tasks. Compared to the text-only frontier model jina-embeddings-v3, performance is lackluster on retrieval tasks. This suggests that multimodal training is hampering text-only performance.
@@ 4.3 Visual Document Retrieval
| Task |
| Retrieval - nDCG@5 [%] | ||||||||||||||||||
| ArxivQ | 64.92 | 58.78 | 66.14 | 25.40 | 30.44 | 43.2 | ||||||||||||
| DocQ | 24.64 | 21.16 | 24.81 | 11.90 | 23.82 | 30.3 | ||||||||||||
| InfoQ | 57.90 | 56.62 | 56.63 | 35.50 | 59.87 | 64.1 | ||||||||||||
| TabF | 45.91 | 37.28 | 43.18 | 20.20 | 70.68 | 58.1 | ||||||||||||
| TATQ | 30.25 | 12.09 | 25.79 | 3.30 | 20.00 | 26.2 | ||||||||||||
| Shift | 34.07 | 28.41 | 31.54 | 3.80 | 30.79 | 18.7 | ||||||||||||
| AI | 68.07 | 32.95 | 55.64 | 15.20 | 47.90 | 62.5 | ||||||||||||
| Energy | 62.15 | 49.15 | 60.97 | 19.70 | 64.94 | 65.7 | ||||||||||||
| Gov | 68.97 | 37.81 | 55.56 | 21.40 | 58.62 | 66.1 | ||||||||||||
| Health | 69.05 | 39.50 | 57.37 | 20.80 | 58.43 | 79.1 | ||||||||||||
| Average | 52.65 | 37.37 | 47.76 | 17.72 | 46.55 | 51.4 | ||||||||||||
Visual document retrieval challenges vision-language models to capture fine-grained information from images, including embedded text, and accurately match these with text queries or documents. This task is particularly challenging as it deals with high-resolution document images with complex content, which differs significantly from processing typical image-caption datasets. As shown in Table 4, jina-clip-v2 outperforms other state-of-the-art CLIP models on the ViDoRe Benchmark for visual document understanding (Faysse et al., 2024), with an average nDCG@5 score of 52.65%, toping jina-clip-v1, nllb-siglip-large and siglip-so400m-patch14-384 (Zhai et al., 2023; Alabdulmohsin et al., 2024). The progression through training stages brings substantial improvements, highlighting the effectiveness of gradually increasing the image resolution. We further investigate how image resolution affects visual document retrieval in Section 5.1.
@@ 4.4 Matryoshka Representation Learning
| CLIP Benchmark | 79.10 | 79.12 | 78.93 | 78.32 | 75.90 | 70.51 | ||||||
| Crossmodal-3600 (Thapliyal et al., 2022) | 81.43 | 82.35 | 82.31 | 81.75 | 78.17 | 72.52 | ||||||
| XTD10 (Aggarwal & Kale, 2020; Rajendran et al., 2016) | 84.87 | 84.85 | 84.60 | 84.32 | 81.80 | 77.85 | ||||||
@@
| CLIP Benchmark | 89.73 | 89.60 | 89.55 | 89.35 | 87.48 | 83.20 | ||||||
| Crossmodal-3600 (Thapliyal et al., 2022) | 83.23 | 83.26 | 83.21 | 82.81 | 80.54 | 75.37 | ||||||
| XTD10 (Aggarwal & Kale, 2020; Rajendran et al., 2016) | 86.03 | 86.02 | 86.02 | 85.84 | 84.37 | 81.02 | ||||||
@@
| EN Classic MTEB (Muennighoff et al., 2023) STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
Table 5 presents the impact of embedding truncation on the crossmodal and text-only tasks. Performance remains highly stable when reducing dimensions from 1024 to 256, with minimal degradation (typically less than 1%) across all evaluation tasks. At 256 dimensions, representing a 75% reduction in embedding size, the model preserves effectively all essential semantic information. Substantial performance degradation is only observed at dimensions 128 and 64.
@@ 5 Analysis
5.1 The Role of Image Resolution in Visual Document Retrieval
@@
Our analysis of performance at various stages of training demonstrates that image resolution plays a critical role in retrieving visually rich documents. To better understand the relationship between image resolution and model performance on such documents, we conduct a targeted experiment.
@@
We checkpointed jina-clip-v2 at the end of Stage 1, when it had only seen images with a resolution of . We conducted four runs, keeping the image resolution to for the first run and increasing to , , and for three additional runs. Each run was trained for 3500 steps using the same visually rich training set and hyperparameters. The models were then evaluated on the ViDoRe benchmark (Faysse et al., 2024), which comprises 10 datasets designed for retrieving visually rich documents with text queries. We plot the results in Figure 2.
@@
Unsurprisingly, increasing image resolution has a positive impact on linking queries to visually rich documents. The most significant improvement occurs when resolution increases from to , with the average nDCG@5 score across 10 benchmarks rising from to . A further increase in resolution to also brings a noticeable gain in performance.
@@
Considering that the model is trained with a patch size of 14, increasing the resolution beyond raises the number of patches quadratically, for example, by a factor of when increasing the resolution to . This large increase in computing costs yields only marginal improvements, with nDCG@5 increasing by just . We consider the optimal image resolution for this model, with a reasonable balance between performance and cost.
5.2 Unified Batch Training vs Multi-Task Learning
@@
Our training approach, as outlined in Section 3, involves two tasks at each stage, both optimized in a multi-task learning framework: a multimodal alignment task for images and text, and a second text-only task. Both use the InfoNCE (Van den Oord et al., 2018) objective. Regardless of whether the loss is computed on image-text or text-text pairs, the same loss function is applied to pairs of vectors. This raises the question: Is there any value in merging the image-text pair and text-text pair data into mixed modality batches and optimizing a single contrastive objective?
@@
We refer to this approach as the Unified Batch Training technique, in contrast to the Multi-Task Learning paradigm. The motivation is two-fold: simplicity of computation, and the possibility of mitigating the modality gap (Liang et al., 2022) by using in-batch negatives across modalities. Compared to multi-task training, the unified batch approach would use a single InfoNCE loss with a shared temperature value, attempting to force all modalities to align within the same embedding space. For this investigation, we also integrated a SimCLR-like self-supervised learning method (Chen et al., 2020) into the contrastive training, aiming to provide more robust image representations. A schematic representation of the unified batch technique is given in Figure 3 in Section A.
@@
Our Stage 1 experiments show that this technique does bring some early improvements, but fails beyond a certain point, at least with regard to crossmodal tasks. This was reflected in the slow decay of the loss temperature (our temperature is a trainable parameter in this case) and its eventual stabilization around . This is in contrast to the multi-task learning approach, where the temperature value decreases more rapidly and keeps decreasing throughout training down to , leading to better performance on crossmodal tasks.
@@
We hypothesize that this limitation arises from the fundamental information asymmetry between the visual and textual modalities (Schrodi et al., 2024). Images are very information-dense, full of small details, while the corresponding textual descriptions are far less detailed (Schrodi et al., 2024; Chen et al., 2023b). In temperature-scaled contrastive learning, the temperature parameter controls the hardness of the loss function with respect to the negative samples, i.e., the level of penalties on the hard negative samples (Wang & Liu, 2021). From an information-theory perspective, the level of penalty on the hard negative text-text samples theoretically differs from that of image-text samples due to the information gap problem. Consequently, as supported by our experiments, enforcing a unified temperature parameter across both modalities is empirically suboptimal.
@@ 6 Conclusion
This work presents an enhanced strategy for training dual-encoder vision-language embedding models for multilingual crossmodal as well as text retrieval and semantic text similarity tasks using contrastive learning. We introduce improvements to the training, namely support for multiple languages, Matryoshka Representation Learning, and visually-rich document understanding. We have used these improvements to train and release a new model, jina-clip-v2, with strong crossmodal and text-only performance on standard benchmarks. Finally, we have presented in this paper analyses of two important considerations for CLIP models going forward: the effects of the modality gap and the role of the image resolution in understanding complex visual inputs.
@@ Appendix A Appendix
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||||
| average | 81.43 | 73.51 | 84.13 | 16.93 | 82.07 | 79.29 | ||||||||||
| ar | 73.56 | 66.22 | 76.89 | 0.19 | 78.92 | 76.94 | ||||||||||
| bn | 63.78 | 50.19 | 68.11 | 0.11 | 75.19 | 74.19 | ||||||||||
| da | 85.39 | 76.53 | 87.97 | 15.39 | 87.14 | 86.64 | ||||||||||
| de | 91.25 | 84.64 | 92.42 | 37.42 | 89.56 | 87.25 | ||||||||||
| el | 75.03 | 68.53 | 77.92 | 0.56 | 77.83 | 71.97 | ||||||||||
| en | 75.83 | 69.78 | 77.78 | 76.17 | 73.11 | 72.22 | ||||||||||
| es | 83.64 | 78.28 | 86.28 | 47.36 | 82.64 | 79.97 | ||||||||||
| fi | 82.83 | 75.28 | 85.89 | 5.03 | 86.42 | 81.44 | ||||||||||
| fr | 88.78 | 83.50 | 90.67 | 62.33 | 87.86 | 85.58 | ||||||||||
| hi | 55.25 | 42.03 | 59.47 | 0.11 | 60.31 | 58.08 | ||||||||||
| id | 84.22 | 73.69 | 86.94 | 13.14 | 86.31 | 84.17 | ||||||||||
| it | 88.33 | 81.19 | 89.64 | 33.39 | 85.94 | 82.67 | ||||||||||
| ja | 87.03 | 77.28 | 90.03 | 3.75 | 86.06 | 83.14 | ||||||||||
| ko | 78.81 | 71.81 | 83.22 | 0.33 | 78.75 | 75.47 | ||||||||||
| nl | 82.56 | 75.72 | 84.47 | 27.69 | 81.69 | 78.86 | ||||||||||
| no | 81.08 | 71.97 | 83.08 | 15.61 | 82.69 | 79.97 | ||||||||||
| pl | 84.00 | 79.33 | 86.50 | 7.39 | 82.72 | 78.61 | ||||||||||
| pt | 82.42 | 76.17 | 85.19 | 31.97 | 82.69 | 79.44 | ||||||||||
| ro | 89.36 | 82.92 | 92.22 | 17.89 | 90.03 | 86.17 | ||||||||||
| ru | 88.97 | 82.97 | 91.11 | 2.19 | 86.44 | 83.78 | ||||||||||
| sv | 78.06 | 71.33 | 80.56 | 15.22 | 79.33 | 76.17 | ||||||||||
| th | 81.61 | 70.92 | 85.08 | 1.89 | 81.14 | 78.83 | ||||||||||
| tr | 81.31 | 75.31 | 84.53 | 4.86 | 83.47 | 81.00 | ||||||||||
| uk | 88.56 | 82.28 | 89.89 | 0.94 | 85.44 | 81.89 | ||||||||||
| vi | 86.64 | 76.31 | 89.06 | 3.17 | 85.56 | 82.56 | ||||||||||
| zh | 78.97 | 66.97 | 82.50 | 2.06 | 76.56 | 74.64 | ||||||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||||
| average | 83.23 | 79.64 | 84.12 | 19.82 | 80.16 | 76.56 | ||||||||||
| ar | 76.25 | 72.14 | 76.67 | 0.31 | 75.86 | 73.69 | ||||||||||
| bn | 69.00 | 63.78 | 70.22 | 0.11 | 75.58 | 73.61 | ||||||||||
| da | 88.53 | 84.42 | 88.86 | 17.25 | 86.53 | 84.69 | ||||||||||
| de | 92.47 | 88.42 | 92.47 | 39.75 | 87.50 | 84.44 | ||||||||||
| el | 73.33 | 73.22 | 75.61 | 0.64 | 74.81 | 68.89 | ||||||||||
| en | 78.58 | 74.33 | 79.61 | 79.06 | 70.81 | 69.08 | ||||||||||
| es | 86.28 | 81.19 | 87.36 | 49.14 | 81.19 | 77.00 | ||||||||||
| fi | 84.19 | 81.17 | 85.64 | 7.22 | 84.25 | 78.75 | ||||||||||
| fr | 90.89 | 86.56 | 91.00 | 63.14 | 86.64 | 83.64 | ||||||||||
| hi | 61.64 | 54.39 | 61.56 | 0.08 | 61.89 | 59.83 | ||||||||||
| id | 86.31 | 83.64 | 87.64 | 16.75 | 84.33 | 81.89 | ||||||||||
| it | 90.17 | 84.31 | 90.53 | 36.53 | 83.50 | 78.64 | ||||||||||
| ja | 88.50 | 85.53 | 89.44 | 7.69 | 84.03 | 81.14 | ||||||||||
| ko | 81.42 | 79.22 | 83.00 | 0.53 | 76.75 | 74.17 | ||||||||||
| nl | 82.47 | 77.94 | 83.33 | 29.28 | 79.28 | 74.33 | ||||||||||
| no | 83.75 | 77.22 | 84.42 | 18.53 | 82.08 | 77.03 | ||||||||||
| pl | 84.61 | 83.72 | 85.69 | 10.00 | 79.58 | 75.39 | ||||||||||
| pt | 83.94 | 79.31 | 84.06 | 33.33 | 79.39 | 74.89 | ||||||||||
| ro | 91.31 | 87.39 | 92.08 | 18.83 | 88.83 | 83.58 | ||||||||||
| ru | 90.64 | 87.75 | 90.89 | 3.33 | 84.64 | 80.61 | ||||||||||
| sv | 78.28 | 77.53 | 80.11 | 16.50 | 76.58 | 72.31 | ||||||||||
| th | 81.94 | 80.22 | 84.11 | 2.17 | 79.56 | 76.28 | ||||||||||
| tr | 82.67 | 80.36 | 83.81 | 6.94 | 80.36 | 77.69 | ||||||||||
| uk | 89.53 | 86.42 | 89.97 | 1.58 | 83.22 | 78.67 | ||||||||||
| vi | 88.06 | 85.64 | 88.58 | 6.06 | 83.14 | 80.03 | ||||||||||
| zh | 79.22 | 74.94 | 80.42 | 4.67 | 73.83 | 70.22 | ||||||||||